Is the function f :[3, 4] ➡R defined by f(x) = x^2 -x is a monotonic in its domain

1
Expert's answer
2021-11-01T17:50:46-0400
"f(x)=x^2-x, x\in[3, 4]"

"f'(x)=(x^2-x)'=2x-1"

If "x\in[3, 4]," then "f'(x)>0" and "f(x)" is (strictly) increasing on "[3, 4]."

Therefore the function "f(x)=x^2-x, x\in [3,4]" is a monotonic in its domain.